A marathon runner runs 6 miles south and then 8 miles east. How far is the runner from the starting point? (Hint: use Pythagorean Theorem)

Answer by Tobiasz(54)

ok
6*(2) + 8*(2) = c*(2)
36 + 64 = C*(2)
100 = c*(2)
^100 = ^c
10 = c